# Empowering Biomedical Discovery with AI Agents

### v1

@@ 4 Levels of autonomy in AI agents
Level 0, denoted as no AI agent, using ML models as tools. This level aligns with the prevailing approach of using interactive and foundation learning models (Section 2). At this level, ML models do not independently formulate testable and falsifiable statements [77] as hypotheses. Instead, model outputs help scientists to form precise hypotheses. For example, a study employed AlphaFold-Multimer to predict interactions of DONSON, a protein with limited understanding, leading to a hypothesis about its functions [78]. Level 1, termed AI agent as a research assistant, features scientists setting hypotheses, specifying necessary tasks to achieve objectives, and assigning specific functions to agents. These agents work with a restricted range of tools and multi-modal data to execute these tasks. For instance, ChemCrow [3] combines chain-of-thought reasoning [79] with ML tools to support tasks in organic chemistry, identifying and summarizing literature to inform experiments. In another example, AutoBa [80] automates multi-omic analyses. These two agents are designed for narrow scientific domains; ChemCrow and AutoBa optimize and execute actions to complete tasks that are designed and predefined by scientists. Level 1 agents [81, 82, 80, 3] formulate simple hypotheses inferred from existing knowledge and utilize a limited set of tools, lacking the capacity for a major scientific discovery necessary to achieve Level 2 autonomy.
@@
At Level 2, AI agent as a collaborator, the role of AI expands as scientists and agents collaboratively refine hypotheses. Agents undertake tasks critical for hypothesis testing, using a wider array of ML and experimental tools for scientific discovery. However, their capability to understand scientific phenomena and generate innovative hypotheses remains constrained, highlighting a linear progression from existing studies. The transition to Level 3, or AI agent as a scientist, marks a major evolution, with agents capable of developing and extrapolating hypotheses beyond the scope of prior research, synthesizing concepts beyond summarizing findings, and establishing concise, informative and clear conceptual links between findings that cannot be inferred from literature alone, eventually yielding a new scientific understanding. While multiple Level 1 agents exist across various scientific fields, Levels 2 and 3 agents have yet to be realized.
@@
As the level of autonomy increases, so does the potential for misuse and the risk of scientists developing an overreliance on agents. While agents have the potential to enhance scientific integrity, there are concerns regarding their use in identifying hazardous substances or controlled substances [84]. Responsible development of agents requires developing preventive measures; further discussion on risks is in [85, 86] and Section 6. The responsible deployment of agents must account for the risk of overreliance, particularly in light of evidence that LLMs can produce convincing but misleading claims and spread misinformation. The risks will likely increase as agents undertake more autonomous research activities. Agents must be subjected to the same scrutiny as scientists, including reproducibility and rigorous peer review of agentic research.
4.1 Illustration of AI agents in genetics
@@
Research in human genetics seeks to understand the impact of DNA sequence variation on human traits. LLM-based agents operating at Level 1 would perform specific tasks relevant to genetic studies. For instance, in a genome-wide association study (GWAS), a Level 1 agent can write bioinformatics code to process raw genotype data to (1) execute quality control measures, such as the removal of single-nucleotide polymorphisms (SNPs) missing in many individuals or control for population stratification [87], (2) estimate ungenotyped SNPs through imputation, and (3) conduct the appropriate statistical analyses to identify relevant SNPs, taking into account the false discovery rate [88]. Following the analysis, the Level 1 agent reviews and reports findings, including any filtered SNPs and rationales for their exclusion.
@@
Instead of executing narrow tasks following human instruction, a Level 2 agent identifies and then executes tasks on its own in order to refine a hypothesis initially given by the scientist. For example, it may explore the effectiveness of drugs for a patient subgroup within complex diseases, where genetic underpinnings can influence drug response [89]. Given a hypothesis that a particular drug is effective in a subset of patients with idiopathic or genetic generalized epilepsy (GGE)—a condition with a robust genetic causality [90]—a Level 2 agent would synthesize genetic information from GWAS meta-analyses [91], such as the UK Biobank [92], targeted sequencing studies [93], and knowledge bases like Genes4Epilepsy [94]. The agent identifies GGE subtypes and causal genes by analyzing patient genetic data, predicting which subgroups might benefit from the drug based on genetic markers. It would then conduct in vitro functional studies to confirm these predictions, ultimately presenting evidence on how the drug could benefit GGE patient subpopulations by synthesizing concepts beyond summarizing findings.
@@
4.2 Illustration of AI agents in cell biology
@@
Cells are fundamental units of study in cell biology. Advances in single-cell omics, super-resolution microscopy, and gene editing have generated datasets on normal and perturbed cells, covering areas such as multi-omics [97, 98, 99], cell viability [100], morphology [101, 102], cryo-electron microscopy and tomography [103, 104], and multiplexed spatial proteomics [105, 106]. This proliferation of data has spurred interest in in silico cell modeling [107] with the overall goal of identifying new mechanistic understanding and adaptive cellular function.
@@
ML tools have been instrumental in analyzing data across these cellular modalities, albeit as Level 0 agents these tools lack autonomous research capabilities. At Level 1, agents integrate specialized Level 0 models designed for narrow tasks such as cell type annotation to assist in hypothesis testing and answering scientific questions. These agents actively engage in research, synthesizing literature and predicting cellular responses using integrated models. For example, to investigate a new compound’s mechanism of action (MoA) [99], a Level 1 agent predicts its effects in various cellular contexts [108]. These predictions inform experimental design, such as transcriptomic and proteomic screening. The agent also retrieves and refines experimental protocols for execution on platforms [109]. After data generation, the agent applies predefined bioinformatics pipelines to extract information from multi-modal data before scientists interpret the results.
@@
Level 2 agents not only execute predefined tasks but also generate hypotheses on cellular functions and compound mechanisms. They autonomously refine tasks to support scientific reasoning, enabling more efficient exploration of complex phenotypes like drug resistance through combinatorial perturbations [110]. By managing the experimental cycle and continuously updating their in silico tools, Level 2 agents aim to minimize experimental redundancy and focus on key variables of resistance. However, efficient probing of the intractable combinatorial space of genes and contexts is still a considerable challenge. In light of this, level 2 agents would propose and optimize experiments for a comprehensive yet cost-effective evaluation of hypotheses through knowledge bases, managing the entire experimental cycle from protocol retrieval and improvement to execution, based on iterative feedback from scientists [56]. Automatic data analysis occurs at every experimental stage, with the agent refining hypotheses, adjusting plans, and continuously updating its ML tools based on a synthesis of predictive content, uncertainty, and newly acquired data.
@@
Level 3 agent systems will implement hybrid cell models as digital-experimental simulators of cellular responses to any type of perturbation. Agents in the hybrid cell model system combine AI tools (digital agents) with high-throughput platforms (experimental agents). Digital agents, such as LLM-based agents, autonomously identify knowledge gaps through literature synthesis and handle any perturbagen (extrinsic events such as gene knockouts, compounds, cell-cell interactions; intrinsic events such as cell cycle) based on criteria such as data volume, biological relevance, and clinical needs. Experimental agents generate high-throughput transcriptomics [111] and proteomics measurements [112, 113] with a spatial resolution for perturbations, perturbation screens including in vivo sequencing, and mass spectrometry platforms [113], as well as barcoded spatial proteomics [112, 114]. Optimization of experimental protocols [115, 101, 116] shifts the role of scientists from performing minute operational tasks to managing hybrid cell models.
4.3 Illustration of AI agents in chemical biology
@@
A major focus of chemical biology scientists is on molecular interactions within cells to manipulate biological systems at molecular and cellular levels. An ideal AI agent would be able to analyze any molecular interaction, help design new drugs, and provide more valuable chemical probes for biological systems.
Despite considerable advances in applying ML to chemical biology, progress toward an AI agent proficient in the scientific method within this field remains limited. Existing predictive models for protein structure prediction [39], molecular docking [117], and generative models for protein design [118, 119] are ML tools that can be wrapped within agents through ‘function calling’ and APIs (Section 5.2). Such agents would be categorized as Level 0, where a scientist oversees all activities. A Level 1 agent, capable of studying a specific protein target, integrates ML tools, such as AlphaFold for structure prediction [39] and neural networks for screening chemical libraries [120] to find candidate chemical compounds that might bind the target. This agent would possess elementary reasoning abilities by probing relevant literature and designing experiments within a specific domain [2], although it would require human guidance for complex inquiries and error correction. The scientist will need to guide the agent to assist with more complex questions and for error checking. For example, a Level 1 agent tasked with binder design for a well-known target would design very similar drug derivatives to existing binders, for which the scientist would have to guide the agent with many different drug scaffolds. However, for a more challenging target, it would provide nonsensical binders.
A Level 2 agent would surpass Level 1 by identifying novel research avenues, such as designing new binders by leveraging trends in related targets; a method demonstrated in materials science for discovering new materials [121]. This agent would employ advanced reasoning extracting literature with more abstract connections to the research task, such as identifying scaffolds that bind to similar pockets and adapting them for the target. In addition, the agent has broader capabilities in planning and designing experiments than a Level 1 agent with expertise in more domains such as retro-synthesis, crystallography, bioassays, and directing robotic arms to execute these experiments. It can also reason from the experiments’ results to identify novel discoveries and propose follow-up experiments. Thus, the agent demonstrates a more robust understanding of chemical biology across hypotheses, experiments, and reasoning. The scientist collaborates with the agent by providing feedback on the quality of designed inhibitors and guiding the agent’s decisions.
A Level 3 agent is capable of studying all types of molecular interactions in a cell. This agent would work alongside human scientists to explore research questions, proposing de novo binders for an undruggable target [122] or a poorly studied target. Unlike the Level 2 agent’s use of well-established protocols, a Level 3 agent unlocks experimental capabilities at spatial and temporal scales that are not currently accessible to experimentation, or design of in situ experiments to study molecular interactions. We can envision the agent proposing candidate molecules, synthesizing molecules with more complex pathways, and designing and executing assays to test efficacy. Finally, it contextualizes results with literature to understand what chemical interactions would occur and why they would be prevalent.
5 Roadmap for building AI agents
@@
AI agents can be obtained through LLM-based systems augmented with different modules [4, 58, 59] that implement functionalities and accomplish tasks detailed in Sections 3 and 4. Here, we describe these modules (Figure 4), focusing on perception, interaction, memory, and reasoning that are necessary for AI agents to engage with human and experimental environments and to make decisions. Interactions between an agent and its environment are characterized by two elements: the agent’s perception of its surroundings and its subsequent engagement with them. Perception modules enable the agent to interpret and assimilate information from various data modalities. Then, learning and memory allow agents to interact with an environment and complete tasks, by acquiring new knowledge and retrieving previously learned one. Finally, Reasoning modules process information and execute action plans. Figure 5e illustrates an agent system that uses perception, interaction, memory, and reasoning modules for research in cell biology [123].
5.1 Perception modules
@@
Perception modules equip LLM-based agents with the capability to understand and interact with elements in the environment in which they operate, such as biological workflows and human users. For perception, agents need to integrate abilities to receive feedback from multiple sources: scientists [50], the environment [63], and other AI agents [14, 124]. This requires accommodating a diverse array of modalities. These include: text descriptions [7]; images from light and (cryo-)electron microscopy to assess cellular processes across many conditions simultaneously [103, 104, 125]; videos from live imaging to assess developmental processes or animal behaviors across time [126, 127, 128]; longitudinal biosensor readouts and genomics profiles of cells [129]; mass spectrometry-based proteomics to decipher protein homeostasis [130, 25]; and miniaturized platforms for conducting biochemical assays and 3D culture systems that mimic the physiological context of organ systems [109].
@@
AI agents can take different approaches to interacting with environments. The most direct one involves using natural language, which represents a common perception modality for LLM-based agents. Other techniques involve using multi-modal perception modules, where agents process multi-modal data streams from the environment or align multi-modal inputs with text-based LLMs.
@@
Multi-modal perception modules. Agents align LLMs with other data types to consider data modalities beyond natural language. This approach helps agents better model the changing environment in which the agent acts and dynamically adjust its outputs to new situations, such as evolved biological states in a virtual cell model. The alignment process involves two main strategies: textual translation and representation alignment. Textual translation converts inputs into a textual format, such as transforming data from robotics into textual descriptions that log environmental states [10]. Alternatively, through representation alignment, data from different modalities are analyzed by modality-specific models to generate representations, such as using the visual encoder from CLIP [131] for visual information processing. These representations are then aligned with LLM textual representations [132, 133, 134, 125, 135] through instruction tuning [125], enabling agents powered by LLMs to perceive and interpret multi-modal data. An alternative to alignment involves allowing the agents to receive input expressed in different modalities [8, 136]. For instance, Fuyu [136] uses a decoder-only transformer architecture to process image patches and text tokens jointly. Similarly, Gemini [8] is engineered to handle visual, audio, and text inputs within a single model. Once perception modules are implemented for agents to receive inputs from the environment, modules for interaction (Section 5.2) and reasoning (Section 5.4) follow to process the inputs and interact externally.
5.2 Interaction modules
@@
Beyond conversational modules, in biological research, scientists use ML-based and other tools, explore datasets through graphical user interfaces (GUIs) to analyze and visualize data, and engage with physical equipment and wet lab experimental platforms. Chat-optimized LLM-based agents thus need interaction capabilities to communicate and collaborate with scientists, other AI agents, and tools to function beyond a simple chatbot. Agents must incorporate essential interaction modules to interact with elements in the environment. These include agent-human interaction to support communication with scientists and following human instruction [137, 138], multi-agent interaction for collaboration among agents, and tool-use action to access ML tools and experimental platforms.
@@
These interaction abilities of LLMs, when combined with interactive ‘function calling’ (i.e., LLM requesting for tasks to be completed), can act as an intermediary between scientists and the agent’s interface, as well as between scientists and various functional items (e.g., tools, other agents, humans). This allows scientists not to search where and how to accomplish tasks but to simply express their intentions in their language. At the same time, the advantages of functional items are fully preserved because agents can interact with tools and use them to provide feedback.
@@
Multi-agent interaction. Multi-agent interactions support solving complex goals that agents could not complete if they operated independently of each other. In such interdisciplinary systems, agents that could specialize in different biological domains, each with distinct capabilities, engage in interactions through various communication means. Language has emerged as the predominant medium for multi-agent interactions due to the ability of agents to communicate with humans linguistically [5, 140, 14, 124, 74]. An instance of this is Generative agents [63], which create interactive environments where agents mimic human behavior and interact using natural language. Different strategies are used for multi-agent interaction, including cooperation [141, 142, 143] and negotiation [74, 144, 145]. For example, MetaGPT [142] applies standardized operating procedures from human teamwork to define tasks and agent responsibilities.
@@
5.3 Memory and learning modules
@@
5.4 Reasoning modules
Biological research involves a multidisciplinary and multistage process that integrates the expertise of scientists from various disciplines. These scientists formulate hypotheses, design experimental setups based on these hypotheses, interpret the results, and plan the next steps. Central to this process is human reasoning, an ability that conventional ML models find challenging to replicate. The integration of reasoning capabilities in AI agents holds the potential to expedite biological research by assisting in several of these critical steps. Reasoning improves agents’ capabilities to plan experiments, make decisions on biological hypotheses, and re-solve competing candidate biological mechanisms. Reasoning modules can be implemented using direct prompting [170] and few-shot in-context learning [79]. Additionally, agents can use planner models [171, 172] and action models [167]. We classify reasoning modules into two categories: direct reasoning and reasoning with feedback. The classification depends on whether the agent adjusts its plan in response to experimental or human feedback.

### v2

@@ Introduction
Level 0, denoted as no AI agent, uses ML models as tools coordinated by interactive and foundation learning models. At this level, ML models do not independently formulate testable and falsifiable statements [78] as hypotheses. Instead, model outputs help scientists to form precise hypotheses. For example, a study employed AlphaFold-Multimer to predict interactions of DONSON, a protein with limited understanding, leading to a hypothesis about its functions [79]. Level 1, termed AI agent as a research assistant, features scientists setting hypotheses, specifying necessary tasks to achieve objectives, and assigning specific functions to agents. These agents work with a restricted range of tools and multimodal data to execute these tasks. For instance, ChemCrow [2] combines chain-of-thought reasoning [80] with ML tools to support tasks in organic chemistry, identifying and summarizing literature to inform experiments. In another example, AutoBa [81] automates multi-omic analyses. These two agents are designed for narrow scientific domains; ChemCrow and AutoBa optimize and execute actions to complete tasks that are designed and predefined by scientists. Level 1 agents [82, 83, 81, 2] formulate simple hypotheses inferred from existing knowledge and utilize a limited set of tools, lacking the capacity necessary to achieve Level 2 autonomy.
@@
At Level 2, AI agent as a collaborator, the role of AI expands as scientists and agents collaboratively refine hypotheses. Agents undertake tasks critical for hypothesis testing, using a wider array of ML and experimental tools for scientific discovery. [76] However, their capability to understand scientific phenomena and generate innovative hypotheses remains constrained, highlighting a linear progression from existing studies. The transition to Level 3, or AI agent as a scientist, marks a major evolution, with agents capable of developing and extrapolating hypotheses beyond the scope of prior research, synthesizing concepts beyond summarizing findings and establishing concise, informative, and clear conceptual links between findings that cannot be inferred from literature alone, eventually yielding a new scientific understanding. While multiple Level 1 agents exist across various scientific fields, Levels 2 and 3 agents have yet to be realized.
@@
As the level of autonomy increases, so does the potential for misuse and the risk of scientists developing an overreliance on agents. While agents have the potential to enhance scientific integrity, there are concerns regarding their use in identifying hazardous substances or controlled substances [85]. Responsible development of agents requires developing preventive measures [86, 87]. The responsible deployment of agents must account for the risk of overreliance, particularly in light of evidence that LLMs can produce convincing but misleading claims and spread misinformation. The risks will likely increase as agents undertake more autonomous research activities. Agents must be scrutinized as scientists, including reproducibility and rigorous peer review of agentic research. We illustrate these definitions of levels by giving examples of progression between the levels in genetics, cell biology, and chemical biology. We selected these areas because of the availability of large datasets that have recently driven the development and application of ML models. We describe the challenges and limitations that biomedical AI agents may present for an enhanced understanding and progression through levels of autonomy.
Illustration of AI agents in genetics
@@
Research in human genetics seeks to understand the impact of DNA sequence variation on human traits. LLM-based agents operating at Level 1 would perform specific tasks relevant to genetic studies. For instance, in a genome-wide association study (GWAS), a Level 1 agent can write bioinformatics code to process genotype data to (1) execute quality control measures, such as the removal of single-nucleotide polymorphisms (SNPs) missing in many individuals or control for population stratification [88], (2) estimate ungenotyped SNPs through imputation, and (3) conduct the appropriate statistical analyses to identify relevant SNPs, taking into account the false discovery rate [89]. Following the analysis, the Level 1 agent reviews and reports findings, including any filtered SNPs and rationales for their exclusion.
@@
Instead of executing narrow tasks following human instruction, a Level 2 agent identifies and executes tasks independently to refine a hypothesis initially given by the scientist. For example, it may explore the effectiveness of drugs for a patient subgroup within complex diseases, where genetic underpinnings can influence drug response [90]. Given a hypothesis that a particular drug is effective in a subset of patients with idiopathic or genetic generalized epilepsy (GGE)—a condition with a robust genetic causality [91]—a Level 2 agent would synthesize genetic information from GWAS meta-analyses [92], such as the UK Biobank [93], targeted sequencing studies [94], and knowledge bases like Genes4Epilepsy [95]. The agent identifies GGE subtypes and causal genes by analyzing patient genetic data, predicting which subgroups might benefit from the drug based on genetic markers. It would then conduct in vitro functional studies to confirm these predictions, ultimately presenting evidence on how the drug could benefit GGE patient subpopulations by synthesizing concepts beyond summarizing findings.
@@
Illustration of AI agents in cell biology
@@
Cells are fundamental units of study in cell biology. Advances in single-cell omics, super-resolution microscopy, and gene editing have generated datasets on normal and perturbed cells, covering areas such as multi-omics [98, 99, 100], cell viability [101], morphology [102], cryo-electron microscopy and tomography [103, 104], and multiplexed spatial proteomics [105, 106]. This proliferation of data has spurred interest in in silico cell modeling [107].
@@
ML tools have been instrumental in analyzing data across these cellular modalities, but as Level 0 agents, they lack autonomous research capabilities. At Level 1, agents integrate specialized Level 0 models to assist in hypothesis testing. These agents actively assist scientists in developing hypotheses by synthesizing literature and predicting cellular responses using integrated models. For example, to help investigate the resistance mechanism of a compound, Level 1 agents predict its effects in various cellular contexts [108]. These predictions also inform experimental design, such as spatial transcriptomic [109] and proteomic [110, 111] screening. Agents then retrieve and refine experimental protocols for execution on platforms [112] and apply predefined bioinformatics pipelines, as instructed by scientists.
@@
Level 2 agents execute predefined tasks and generate hypotheses on cellular functions and responses. They autonomously define and refine tasks to support scientific reasoning, enabling practical exploration of complex phenotypes like drug resistance. By managing the experimental cycle and continuously updating their in silico tools, Level 2 agents actively optimize experiments to focus on key variables of resistance based on a synthesis of predictive content, uncertainty, and newly acquired data, with iterative feedback from scientists [55]. Level 2 agents thus form a prototype for a virtual cell model capable of hypothesis generation, encompassing closed-loop integration of digital and experimental platforms.
@@
Level 3 agents respond to existing challenges and anticipate future directions in cell biology research. They form hybrid virtual cell models, an organic combination of AI tools (digital agents) with high-throughput platforms (experimental agents). Digital agents, such as LLM-based agents, autonomously identify critical knowledge gaps through literature synthesis based on criteria such as data volume, biological relevance, and clinical needs and simulate any perturbagen (extrinsic events such as gene knockouts and overexpression, compounds, cell-cell interactions; intrinsic events such as cell cycle) in any context. Experimental agents not only optimize experimental protocols [113, 102, 114] to enable high-throughput multimodal measurements but also develop transformative technologies to enable probing at unprecedented resolution across space and time across in vitro, ex vivo, and in vivo models, uncovering pioneering insights. The ability of level 3 agents to drive the discovery of novel biological mechanisms and therapeutic strategies shifts the role of scientists from performing operational tasks to ideation and managing hybrid cell models.
Illustration of AI agents in chemical biology
@@
A major focus for chemical biology is understanding molecular interactions within cells to manipulate biological systems at molecular and cellular levels. An AI agent could analyze any molecular interaction, help design new drugs, and provide more valuable chemical probes for biological systems.
Despite considerable advances in applying ML to chemical biology, current approaches fall in Level 0. Scientists oversee all activities by integrating ML tools for structure prediction, docking, chemical synthesis, and molecular generation. At Level 1 the agent has elementary reasoning of chemical biology and can execute simple tasks autonomously such as running ML tools, or designing experiments for a given objective. However, due to limited reasoning capabilities, the agent may fail to explain more complex concepts, such as how the dynamics of molecules may influence the effects of drugs on binders or explore novel molecular scaffolds. For a level 2, the long-term objective is its function as a collaborator for scientists through excelling at tasks that are explicit continuations of existing scientific research, such as improving the efficiency of chemical probes, autonomously designing and testing de novo enzymes, or designing new binders by leveraging trends in related targets. Level 2 AI agents have deeper expertise in more domains, such as retro-synthesis, crystallography, bioassays, and directing robotic arms to conduct research.
The goal of a Level 3 agent in chemical biology is the ability to study all types of molecular interactions in a cell. This agent would work alongside human scientists to explore research questions that are challenging for the field, such as binder design for undruggable targets [115], significantly improving specificity and efficiency of in vivo bioorthogonal reactions, or developing new chemical probes that can access new spatial and temporal scales. Unlike the Level 2 agent’s use of well-established protocols, a Level 3 agent aims to unlock experimental capabilities that are not currently accessible. For example, AI agents could be tasked to probe molecular dynamics at longer timescales than what is currently accessible. At this level, agents have a thorough understanding of existing literature and work alongside scientists to unlock new fields of chemical biology.
Roadmap for building AI agents
@@
An AI agent is built as a compound system that consists of modules [3, 57, 58] each implementing a distinct functionality. Here, we describe these modules (Figure 4), focusing on perception, interaction, memory, and reasoning modules necessary for AI agents to interact with humans and engage with experimental environments. Interactions between the agent and its environment are characterized by two elements: the agent’s perception of its surroundings and its subsequent engagement with them. Perception modules enable the agent to interpret and assimilate information from various data modalities. Then, learning and memory allow agents to interact with an environment and complete tasks, by acquiring new knowledge and retrieving previously learned one. Finally, the reasoning module processes information and executes action plans. Using a published study as an example [116], Figure 5e illustrates a hypothetical AI agent that examines the molecular mechanisms of selective removal of mitochondrial DNA mutants in the Drosophila female germline through perception, interaction, memory, and reasoning modules.
The division of research into smaller tasks handled by AI agents presents an intriguing approach, building on the success of modular and sequential bioinformatics workflows like Snakemake and Docker. Unlike these workflows, which are often static and require manual updates and reconfiguration to handle new tasks or integrate new tools, AI agents are dynamic and operate in a personalized, user-specific, and context-appropriate manner. They can learn to use new tools and adjust their workflows based on the specific instructions and needs of the scientist. Further, the adaptive allocation of tasks by AI agents can be helpful in automatically incorporating new tools and restructuring existing pipelines, much like a human researcher would. For example, AI agents could experiment with and create new protocols beyond the currently established methods in integrating multimodal omics data. For instance, while established protocols for integrating multi-modal, such as scRNA-seq with scATAC-seq or spatial data, exist, AI agents could develop new pipelines for multi-modal integrations beyond the three modalities, or multi-scale integrations such as atlas-scale single-cell and bulk RNA-seq data, or normal and disease state data from cell lines, organoids, and patient samples, based on their initial attempts.
Perception modules
@@
Perception modules equip LLM-based agents with the capability to understand and interact with elements in the environment in which they operate, such as biological workflows and human users. For perception, agents need to integrate abilities to receive feedback from multiple sources: scientists [49], the environment [62], and other AI agents [117, 13]. This requires accommodating a diverse array of modalities. These include text descriptions [6]; images from light and (cryo-)electron microscopy to assess cellular processes across many conditions simultaneously [103, 104, 118]; videos from live imaging to assess developmental processes or animal behaviors across time [119]; longitudinal biosensor readouts and genomics profiles of cells [120]; mass spectrometry-based proteomics to decipher protein homeostasis [121, 24]; and miniaturized platforms for conducting biochemical assays and 3D culture systems that mimic the physiological context of organ systems [112].
@@
AI agents can take different approaches to interacting with environments. The most direct one involves using natural language, which represents a common perception modality for LLM-based agents. Other techniques involve multi-modal perception modules, where agents process multi-modal data streams from the environment or align multi-modal inputs with text-based LLMs.
@@
Multi-modal perception modules. Agents align LLMs with other data types to consider data modalities beyond natural language. This approach helps agents better model the changing environment in which the agent acts and dynamically adjust its outputs to new situations, such as evolved biological states in a virtual cell model. The alignment process involves two main strategies: textual translation and representation alignment. Textual translation converts inputs into a textual format, such as transforming data from robotics into textual descriptions that log environmental states [9]. For example, when handling readouts from experimental devices, the readouts can be combined with a textual description of their meaning, allowing the LLM to understand the readouts as a new modality. Alternatively, through representation alignment, data from different modalities are analyzed by modality-specific models to generate representations, such as using the visual encoder from CLIP [122] for visual information processing. These representations are then aligned with LLM textual representations through instruction tuning [123, 118], enabling agents powered by LLMs to perceive and interpret multi-modal data. For instance, to make LLM-based agents handle the protein structure data, an additional encoder is required to encode the protein structure data into a representation aligned with the LLMs’ representation space. This encoder is pre-trained with modality-specific training schemes, and an adaptor is placed between this encoder and LLMs to align the representations of the two modalities. Then, instruction tuning is applied using data containing both modalities to train the adaptor for alignment. An alternative to alignment involves allowing the agents to receive input expressed in different modalities [7, 124]. For instance, Fuyu [124] uses a decoder-only transformer architecture to process image patches and text tokens jointly. Similarly, Gemini [7] is engineered to handle visual, audio, and text inputs within a single model. Once perception modules are implemented for agents to receive inputs from the environment, modules for interaction and reasoning follow to process the inputs and interact externally. Training agents with strong perception abilities on biomedical data requires extensive, high-quality data pairs that align multiple modalities. However, collecting such data remains challenging. For example, multimodal experimental platforms are non-existent or have low-throughput yields, certain tissues, and cell types are not experimentally available, and a long tail of disease phenotypes has small sample sizes, making data collection infeasible.
Interaction modules
@@
Beyond conversational modules, scientists use ML-based and other tools in biological research, explore datasets through graphical user interfaces (GUIs) to analyze and visualize data, and engage with physical equipment and wet lab experimental platforms. Chat-optimized LLM-based agents thus need interaction capabilities to communicate and collaborate with scientists, other AI agents, and tools to function beyond a simple chatbot. Agents must incorporate essential interaction modules to interact with elements in the environment. These include agent-human interaction to support communication with scientists and following human instruction [125, 126], multi-agent interaction for collaboration among agents, and tool-use action to access ML tools and experimental platforms.
@@
These interaction abilities of LLMs, when combined with interactive ‘function calling’ (i.e., LLM requesting for tasks to be completed), can act as an intermediary between scientists and the agent’s interface, as well as between scientists and various functional items, such as tools and other agents. This approach allows scientists to express their intentions in natural language without needing to search for how and where to accomplish tasks. At the same time, the advantages of functional items are preserved because agents can interact with tools and use them to provide feedback. However, interactive modules trained on general, non-biological domains might not be well-suited for specialized biomedical terminologies, requiring in-domain training on biomedical tools.
@@
Multi-agent interaction. Multi-agent interactions support solving complex goals that agents could not complete if they operated independently. In such interdisciplinary systems, agents that could specialize in different biological domains, each with distinct capabilities, engage in interactions through various communication means. Language has emerged as the predominant medium for multi-agent interactions due to the ability of agents to communicate with humans linguistically [4, 128, 13, 117, 73]. An instance of this is Generative agents [62], which create interactive environments where agents mimic human behavior and interact using natural language. Different strategies are used for multi-agent interaction, including cooperation [129, 130, 131] and negotiation [73, 132, 133]. For example, MetaGPT [130] applies standardized operating procedures from human teamwork to define tasks and agent responsibilities.
@@
Memory and learning modules
@@
Reasoning modules
Biological research involves a multidisciplinary and multistage process that integrates the expertise of scientists from various disciplines. Scientists formulate hypotheses, design experiments based on these hypotheses, interpret the results, and plan the next steps. The integration of reasoning capabilities in AI agents can assist biological research throughout this process. Reasoning improves agents’ capabilities to plan experiments, make decisions on biological hypotheses, and resolve competing candidate biological mechanisms. AI agents that use large language models can implement interactive dialogue systems to explain ML models through natural language conversations. Reasoning modules can be implemented using prompting [158] and few-shot in-context learning [80]. Additionally, agents can use planner models [159, 160] and action models [155]. We classify reasoning modules into two categories: direct reasoning and reasoning with feedback, depending on whether agents adjust their plan in response to experimental or human feedback.